OpenStudy (anonymous):

Write a polynomial function in standard form with real coefficients whose zeros and their multiplicities include those listed. Graph it and write down your observations of it. -3(multiplicity 2 ), 2(multiplicity 3)

OpenStudy (anonymous):

do you know what "multiplicity" means in this context?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

How do I become Moderator?

OpenStudy (anonymous):

it means the factor as that number as an exponent so the factor that has a zero of \(-3\) is \(x+3\) and since the multiplicity is 2, it is \[(x+3)^2\]

OpenStudy (anonymous):

multiplicity is a factor that repeats more than once

OpenStudy (anonymous):

right so one factor will be \((x+3)^2\) and the other will be \((x-2)^3\) giving you \[f(x)=(x+3)^2(x-2)^3\]

OpenStudy (anonymous):

so for -3(multiplicity 2) will it be (X-3) to the second power

OpenStudy (anonymous):

no, not quite

OpenStudy (anonymous):

it will be \((x+3)^2\)

OpenStudy (anonymous):

if \(x=-3\) then \(x+3=0\) so the factor is \((x+3)^2\) not \((x-3)^2\)

OpenStudy (anonymous):

oook